. A mass m =1 kg located at point (3,4) in x-y plane at time t is subjected to a force of 2 N
in the y direction. All numbers are in SI units. The angular acceleration is
a) 0.24 radians s-2
along z direction b) 0.18 radians s-2
along z direction
c) 0.12 radians s-2
along x direction d) 0.32 radians s-2
along z direction

Grade:

4 Answers

Arun
3 years ago
Dear student
 
Question is not clear. Please check and repost the question with an attachment. I will be happy to help you.
Vikas TU
3 years ago
Ganesh
2 years ago
It is simple T divided by mass moment of inertia
In this case torque is 2*3 =6N-m
Mass moment of inertia is 
5*(3^2+4^2) =25 hence angular acceleration  = 6/25 = 0.24. a is the correct choice
Dr Iruwa
one year ago
I want to quickly write about a Powerful Love spell Caster that helped me when I had a misunderstanding with my husband which led to Divorce. I contacted this Powerful spell caster called Dr Mahnlangu help and I followed the step that he asked and within 24 hours my husband called me and was eager for my forgiveness and that he wished nothing more than to have me in his arms forever. I am filled with so much joy and happiness that I found Dr Mahnlangu . I hope you all here will find this testimony of mine and get your husband back in just 24 hours. He did a perfect job for me. I am So Happy and Excited because as i am writing this Testimony, My husband is madly in love with me again. If there is anybody Out there who is in Difficulties and need help should kindly  whataspp : +27787390989,